# Tunneling in multi-site mesoscopic quantum Hall circuits

### v1

@@ IV Effective low-energy Hamiltonian
By integrating out the gapped modes in the equation (25), we express the backscattering Hamiltonian (4) in terms of a single free mode $\Phi_{a}$. This procedure is elaborated in the appendix B, and the final result for the effective Hamiltonian of the four-site QH circuit writes
@@
where $\theta_{u/v}$ are some unimportant phase factors. In the case of common gate voltage, i.e. $N_{j,\rm g}=N_{\rm g}$, the effective reflection coefficients $|r_{u/v}|$ acquire the form
@@
where $\mathcal{C}_{j}$ and $\mathcal{D}_{j}$ are numerical coefficients as a result of integrating out the gapped bosonic modes, and also contain the proportionality factor of the renormalized backscattering strengths $\mathsf{U}_{j}\propto U_{j}$ and $\mathsf{V}_{j}\propto V_{j}$, see appendix B for the details. The symbol $\gamma$ stands for the Euler’s constant. In the following section, we analyze the quantum critical features of the effective Hamiltonian (27).
@@ V Quantum critical behaviors
At the quantum critical point (a point in the parameter space of the gate voltage and barrier transparencies), all backscattering processes interfere destructively to nullify the backscattering terms in the Hamiltonian (27). At this point, the unitary conductance can be reached (see below). We note that since both backscattering terms are relevant, one is required to find proper parameters to nullify both $|r_{u}|$ and $|r_{v}|$. Although there are several parameters in the equation (27) that can be tunned experimentally, in the following we discuss a few representative cases.
First consider the situation where only the leftmost and rightmost QPCs, i.e., the 1st and 5th, are in quasi-ballistics regime with equal reflection amplitudes $\mathsf{U}_{1}=\mathsf{U}_{5}$ and $\mathsf{V}_{1}=\mathsf{V}_{5}$. Since all the inner QPCs are fully open ( $\mathsf{U}_{2,3,4}=0=\mathsf{V}_{2,3,4}$), the effective reflection coefficients take the form $|r_{u}|\propto\cos 4\pi N_{\rm g}$ and $|r_{v}|\propto\cos 8\pi N_{\rm g}$. Due to the doubling of the gate voltage periodicity in these coefficients, they never vanish simultaneously at any gate voltage. This special realization of the device 1 thus does not support a zero-temperature quantum critical point, i.e., only the perturbative regime would exist, and the dominant scaling behaviors of transport and thermodynamic quantities would be identified by comparing the relative strengths of two corrections in equation (27).
Now we discuss the left-right symmetric realization of the device 1, namely the barrier transparencies are chosen such that $\mathcal{C}_{1}\mathsf{U}_{1}=\mathcal{C}_{2}\mathsf{U}_{2}=\mathcal{C}_{4}\mathsf{U}_{4}=\mathcal{C}_{5}\mathsf{U}_{5}$ and $\mathcal{C}_{3}\mathsf{U}_{3}/\mathcal{C}_{1}\mathsf{U}_{1}=\delta_{u}$, and similarly $\mathcal{D}_{1}\mathsf{V}_{1}=\mathcal{D}_{2}\mathsf{V}_{2}=\mathcal{D}_{4}\mathsf{V}_{4}=\mathcal{D}_{5}\mathsf{V}_{5}$ and $\mathcal{D}_{3}\mathsf{V}_{3}/\mathcal{D}_{1}\mathsf{V}_{1}=\delta_{v}$. In this case, we have $|r_{u}|\propto\left[\delta_{u}+2\cos 2\pi N_{\rm g}+2\cos 4\pi N_{\rm g}\right]$ and $|r_{v}|\propto\left[\delta_{v}+2\cos 4\pi N_{\rm g}+2\cos 8\pi N_{\rm g}\right]$. In this case, the parameter space $\left(\delta_{u},\delta_{v},N_{\rm g}\right)$ forming the critical point exists. As an example, for a given value of $\delta_{u}$, the fine tunned gate voltage
@@
and $\delta_{v}$ such that
$\delta^{*}_{v}=3\alpha+\sqrt{9-4\alpha}-1-\alpha^{2}-\alpha\sqrt{9-4\alpha},$ (30)
constitute one of the quantum critical points. As seen from equation (29), the critical regime emerges only for $0<\delta_{u}\leq 9/4$, and for larger $\delta_{u}$ different critical points meet and nullify each other. One can numerically solve the zero condition of reflection coefficients given in equation (28) to get all the features of critical points for general value of barrier transparencies and five gate voltages. In the following, we show that these features can be accessed via transport measurements.
@@ VI Charge current
Here we evaluate the charge current measured at the right QPC (QPC 5th) described by the operator $\hat{I}=-\frac{1}{2\pi}\frac{\partial\Phi_{5}}{\partial t}$. As discussed earlier, all the fields are a function of $\Phi_{a}$ and since the field $\Phi_{a}$ is only the gapless mode, the application of the voltage bias $V\ll E_{C}$ amounts to the dressing such that $\Phi_{a}\to\Phi_{a}-\left(Vt/\sqrt{5}\right)$. Therefore, in the fully ballistic regime, the charge current is given by
@@
as just equivalent to the conductance corresponding to a setup with five quantum resistors (with resistors $R_{q}=2\pi$) in series.
Figure 2: Schematics representation of the setup 1 in asymmetric tunneling regime (see text for the details).
The backscattering correction to the charge current is accounted for by the usual second order perturbation
@@
Since the two backscattering sectors in equation (27) do no produce finite inter-sector interfere current at the lowest order of perturbation theory (due to the imbalance of charge neutrality condition), equation (32) simplifies to
@@
where $\Gamma(y)$ is the usual Gamma function with argument $y$. Precisely at the critical point discussed earlier, the correction given by the second term in equation (35) vanishes, and the validity of above perturbative results extends down to zero energies. We note that the linear response $V\to 0$ correction to the conductance can be obtained, up to a trivial numerical factor, by replacing $V\to T$ in the second part of the equation (35).
Differential conductance in the asymmetric regime
Here we briefly discuss the situation where the rightmost QPC in Fig. 1 is operating in the tunneling regime while the remaining QPCs are fully ballistic or in the quasi-ballistic regime as schematically shown in Fig. 2. We assume that the contact conductance, i.e. $G_{c}=|\gamma_{t}|^{2}/2\pi$, where $|\gamma_{t}|$ is the tunneling amplitude of the right contact, is the smallest energy scale in the system. In this case, the tunneling current can be expressed in terms of the electron Greens function $G^{>}(t)=-i\Big<\Psi_{d}(t)\Psi_{d}^{\dagger}(0)\Big$> of the fourth grain defined at the tunneling position $x=0$, and that of the usual free electron Greens functions [44]. For the grain, the explicit solution of the equation of motion gives the greater Greens function in the form
@@
where $g^{>}(t)=-i\Big<\Psi_{f}(t)\Psi_{f}^{\dagger}(0)\Big$> is the greater Greens function describing the free electrons, and $Z(\omega)$ is the usual frequency dependent environmental function given by [45, 5]
@@
In the case of the voltage $V$ applied to the free edge channel on the right, the tunneling current can be cast into the form  [^1] The charge current takes the standard form $I=|\gamma_{t}|^{2}\int^{\infty}_{-\infty}dt\;e^{iVt}\Big[g^{<}(t)G^{<}(t)-g^{>}(t)G^{>}(t)\Big]$.
@@
In the asymptotic limit of $T\ll E_{C}$, the equation (38) provides weakly temperature-dependent linear response conductance $G(T)=\mathcal{F}G_{c}(T/E_{C})^{8}$, where $\mathcal{F}$ is a constant of the order of unity. The temperature scaling exponent is governed by the four pairs of electron-hole excitations each contributing $T^{2}$ to the conductance, and thus the setup with large number of sites becomes essentially an electronic insulator at low temperature. All the previous treatments can be applied to compute the finite backscattering corrections to equation (38). Inclusion of such backscattering corrections will amount to the addition of linear in $|r_{u/v}|$ terms to the constant $\mathcal{F}$.
Figure 3: (a) Typical realization of two-channel, two-site quantum Hall cricuits. (b) The setup (a) with one of the channels being looped. The red dot represents finite backscatterings, while in the case of the cross symbol, the corresponding QPC is fully opened. Notice that the gates are not explicitly shown (see text for the details).
@@
VII Multi-channel mesoscopic quantum Hall circuits
In the previous sections, we studied the QH circuits where each grain accommodates just a single channel to the left and also to the right of it. It is also possible to engineer a mesoscopic QH circuit having multiple edge channels passing through each metal grain [23]. The fully ballistic limit of such multichannel QH circuits can be understood as follows. A portion of QH circuit outside the metal grain that consists of left and right moving edges can be described by a chiral bosonic field. Therefore, for the setup with $N$ identical grains each hosting $M$ channels, there would be in total $M(N+1)$ bosonic modes describing the corresponding plasmonic excitations. At low energies as compared to the charging energy and also if the inter-grain Coulomb interactions are negligible, out of $M(N+1)$ bosonic modes, $N$ of them would be gapped by the large charging energy and can be integrated out. The corresponding low energy effective Hamiltonian can then be written in terms of the remaining $N_{\rm f}=\left[M(N+1)-N\right]$ free modes. The usual circuit analysis then shows that the conductance of each free channel takes the value $G_{0}=M/[2\pi(N+1)]$.
In the case of $N_{\rm f}>1$, the problem can in principle be mapped to an equivalent problem of a particle moving in a multidimensional periodic potential in the presence of dissipation [48, 49]. The usual perturbative calculations at large transparencies are generally unstable in the sense that as temperature/voltage tends to zero, the perturbative corrections blow up. To explicitly demonstrate the above articulations, in the following, we study a typical multichannel QH circuit as schematically shown in Fig. 3(a). It consists of two islands and a total of six edge channels. For the clarity of presentation, we call the portion of the circuit to the left of the first island as the left part, that between the two grains as the middle part and that right of the second grain as the right part. Each part accommodates two channels (blue and black), here for generality, implemented via six QPCs as shown in Fig. 3(a).
@@ VII Multi-channel mesoscopic quantum Hall circuits
Exotic quantum critical points in multi-channel QH circuits
As long as there is more than one free mode describing the low energy dynamics, there is no relevance of mapping to the boundary sine-Gordon model. Nonetheless, we show below, following Ref. [23], that in the multichannel QH circuits the number of free modes describing the effective Hamiltonian can always be reduced to one by looping the desired number of edge channels. This method allows realizing zero-temperature quantum critical phenomena characterized by general scaling dimension $\eta<1/2$, where $\eta$ depends on the number of grains and that of looped edge channels.
To illustrate the looping method, we consider a simples setup as shown in Fig. 3(b) formed by looping the second channel of Fig. 3(a) back to the grains [51]. The main idea here is that the transport through the left and right QPCs (QPC 1 and 3 respectively) is unaffected by the presence of looped part (on the left and right) of the circuit. Therefore, if both channels in the middle part are fully ballistic, they only provide the impedances to the incoming electrons via left and right QPCs, the whole setup becomes similar to that with a single grain equipped with two QPCs. The crucial difference here is that the number of connecting ballistic channels provides an extra impedance and thus controls the scaling dimension of the electron operators.
Following the equation of motion approach to account for the charging energy part of the Hamiltonian, one can express all bosonic fields in terms of incoming fields. The fields $\Phi_{\rm o}$ describing the two looped parts of the circuit do not directly enter into our problem of transport through QPCs 1 and 3, and have the correlator of the form
@@
with the gate voltage parameters $N_{\rm L,g}=2\pi\left(3N_{\rm 1g}+2N_{\rm 2g}\right)/5$ and $N_{\rm R,g}=2\pi\left(2N_{\rm 1g}+3N_{\rm 2g}\right)/5$. The boundary term (74) having scaling dimensions $2/5$ supports a zero-temperature quantum critical point that can be reached by varying gate voltages and/or backscattering amplitudes. In addition, here the boundary sine-Gordon model can be exploited to uncover all transport and thermodynamic properties. In the case of $N$ grains and $M$ looped channels, see Fig. 3(b) for $N=2$ and $M=1$, the backscattering in the left and right QPCs can be described by a single gapless field, and the leading oder perturbations have the scaling dimension $\nu_{\rm o}=(M+1)/(2M+N+1)$. All these different realizations support distinct zero-temperature quantum critical points, where for $\nu_{\rm o}<1/4$, the higher order backscattering terms also have to be taken into account. Therefore, the channel-looping in multi-site QH circuits primarily brings new physics and provides far-reaching insights on non-Fermi liquid behaviors.
VIII Heating effect in a multi-site QH circuit
In section VI, we also presented our results for the transport properties of the QH circuit 1 in a fully non-equilibrium regime driven by the voltage bias. There we neglected the effects coming from the heating of the grains by injected voltage, namely the Joules heat $J_{V}=V^{2}/2R_{q}$ carried by each incoming chiral edge due to the corresponding voltage drop $V$. Likewise, at a given temperature $T$, the heat current carried by a free chiral edge described by the bosonic field $\Phi^{0}$ takes the form

### v2

@@ IV Effective low-energy Hamiltonian
By integrating out the gapped modes in the equation (25), we express the backscattering Hamiltonian (4) in terms of a single gapless mode $\Phi_{a}$. This procedure is elaborated in appendix B, and the final result for the effective Hamiltonian of the four-site QH circuit writes
@@
where $\theta_{u/v}$ are unimportant phase factors. In the case of common gate voltage, i.e. $N_{j,\rm g}=N_{\rm g}$, the effective reflection coefficients $|r_{u/v}|$ acquire the form
@@
where $\mathcal{C}_{j}$ and $\mathcal{D}_{j}$ are numerical coefficients arising from integrating out the gapped bosonic modes, and also include proportionality factors associated with the renormalized backscattering strengths $\mathsf{U}_{j}\propto U_{j}$ and $\mathsf{V}_{j}\propto V_{j}$; see appendix B for details. The symbol $\gamma$ stands for the Euler’s constant. In the following section, we analyze the quantum-critical properties of the effective Hamiltonian (27).
@@ V Quantum critical behaviors
At the quantum critical point (a point in the parameter space of the gate voltage and barrier transparencies), all backscattering processes interfere destructively to nullify the backscattering terms in the Hamiltonian (27). At this point, the unitary conductance is achieved (see below). We note that since both backscattering terms are relevant, one is required to find proper parameters to nullify both $|r_{u}|$ and $|r_{v}|$. Although there are several parameters in the equation (27) that can be tunned experimentally, we focus below on a few representative cases.
First consider the situation where only the leftmost and rightmost QPCs, i.e., the first and fifth, are in quasi-ballistics regime with equal reflection amplitudes $\mathsf{U}_{1}=\mathsf{U}_{5}$ and $\mathsf{V}_{1}=\mathsf{V}_{5}$. Since all the inner QPCs are fully open ( $\mathsf{U}_{2,3,4}=0=\mathsf{V}_{2,3,4}$), the effective reflection coefficients take the form $|r_{u}|\propto\cos 4\pi N_{\rm g}$ and $|r_{v}|\propto\cos 8\pi N_{\rm g}$. Due to the different gate-voltage periodicities of these coefficients, they never vanish simultaneously for any gate voltage. This special realization of device 1 thus does not support a zero-temperature quantum critical point, i.e., only a perturbative regime exists, and the dominant scaling behaviors of transport and thermodynamic quantities are determined by the relative strengths of two corrections in equation (27).
We now discuss the left-right symmetric realization of device 1, in which the barrier transparencies are chosen such that $\mathcal{C}_{1}\mathsf{U}_{1}=\mathcal{C}_{2}\mathsf{U}_{2}=\mathcal{C}_{4}\mathsf{U}_{4}=\mathcal{C}_{5}\mathsf{U}_{5}$ and $\mathcal{C}_{3}\mathsf{U}_{3}/\mathcal{C}_{1}\mathsf{U}_{1}=\delta_{u}$, and similarly $\mathcal{D}_{1}\mathsf{V}_{1}=\mathcal{D}_{2}\mathsf{V}_{2}=\mathcal{D}_{4}\mathsf{V}_{4}=\mathcal{D}_{5}\mathsf{V}_{5}$ and $\mathcal{D}_{3}\mathsf{V}_{3}/\mathcal{D}_{1}\mathsf{V}_{1}=\delta_{v}$. In this case, we have $|r_{u}|\propto\left[\delta_{u}+2\cos 2\pi N_{\rm g}+2\cos 4\pi N_{\rm g}\right]$ and $|r_{v}|\propto\left[\delta_{v}+2\cos 4\pi N_{\rm g}+2\cos 8\pi N_{\rm g}\right]$. It is therefore possible to find the parameter space $\left(\delta_{u},\delta_{v},N_{\rm g}\right)$ where $|r_{u}|$ and $|r_{v}|$ vanish simultaneously. As an example, for a given value of $\delta_{u}$, the fine-tuned gate voltage
@@
and $\delta_{v}$ chosen such that
$\delta^{*}_{v}=3\delta_{u}+\sqrt{9-4\delta_{u}}-1-\delta_{u}^{2}-\delta_{u}\sqrt{9-4\delta_{u}},$ (30)
together constitute one of the quantum critical points. As seen from equation (29), the critical regime emerges only for $0<\delta_{u}\leq 9/4$, and for larger $\delta_{u}$ different critical points merge and annihilate each other. One can numerically solve the zero condition of the reflection coefficients given in equation (28) to obtain the full structure of the critical points for general barrier transparencies and gate voltages. In the following, we show that these features can be accessed via transport measurements.
Figure 2: Schematic representation of the setup 1 in the asymmetric tunneling regime (see text for details).
@@ VI Charge current
Here we evaluate the charge current measured at the rightmost QPC (QPC 5), described by the operator $\hat{I}=-\frac{1}{2\pi}\frac{\partial\Phi_{5}}{\partial t}$. As discussed earlier, all fields are function of $\Phi_{a}$, and since $\Phi_{a}$ is the only gapless mode, the application of the voltage bias $V\ll E_{C}$ amounts to the dressing such that $\Phi_{a}\to\Phi_{a}-\left(Vt/\sqrt{5}\right)$. Therefore, in the fully ballistic limit, the charge current is given by
@@
which is equivalent to the conductance of five quantum resistors (with resistors $R_{q}=2\pi$) connected in series.
The backscattering correction to the charge current is accounted for by the usual second-order perturbation theory
@@
Since the two backscattering sectors in equation (27) do no produce finite inter-sector interfere contribution at the lowest order (due to the imbalance of charge neutrality condition), equation (32) simplifies to
@@
where $\Gamma(y)$ denote the Gamma function with argument $y$. Precisely at the critical point discussed above, the correction given by the second term in equation (35) vanishes, and the perturbative treatment remains valid down to zero energies. We note that the linear response $V\to 0$ correction to the conductance can be obtained, up to a trivial numerical prefactor, by replacing $V\to T$ in the second part of the equation (35). The backscattering correction in equation (35) can also be expressed as $\sim\left(T^{*}_{u}/V\right)^{8/5}+\left(T^{*}_{u}/V\right)^{2/5}$, where we introduced the crossover scales $T^{*}_{u}\propto|r_{u}|^{5/4}$ and $T^{*}_{v}\propto|r_{v}|^{5}$. In the general case, the validity of equation (35) is restricted to the region $T^{*}_{u,v}\ll(V,T)$. At the quantum critical point, both crossover scales vanish, and a small detuning of the gate voltage from the critical value $N^{*}_{\rm g}$ (29) gives the dominant scaling behavior $T^{*}_{u}\sim\delta N_{\rm g}^{5/4}$, provided that both $|r_{u}|$ and $|r_{v}|$ remain finite in the vicinity of critical point. Here, $\delta N_{\rm g}=N_{\rm g}-N^{*}_{\rm g}$. If, for a given detuning, $T^{*}_{u}=0$, the scaling is much weaker, with $T^{*}_{v}\sim\delta N_{\rm g}^{5}$.
VII Transport properties in the tunneling regime
We now consider the case where the rightmost QPC in Fig. 1 operates in the tunneling regime, while the remaining QPCs are fully ballistic or in the quasi-ballistic regime as schematically shown in Fig. 2. We assume that the contact conductance, i.e. $G_{c}=|\gamma_{t}|^{2}/2\pi$, where $|\gamma_{t}|$ is the tunneling amplitude of the right contact, is the smallest energy scale in the system. In this case, the tunneling current can be expressed in terms of the electron Greens function $G^{>}(t)=-i\Big<\Psi_{d}(t)\Psi_{d}^{\dagger}(0)\Big$> of the fourth grain defined at the tunneling position $x=0$, and that of the usual free electron Greens functions [47]. For the grain, the explicit solution of the equation of motion gives the greater Greens function in the form
@@ VII Transport properties in the tunneling regime
where $g^{>}(t)=-i\Big<\Psi_{f}(t)\Psi_{f}^{\dagger}(0)\Big$> is the greater Greens function describing the free electrons, and $Z(\omega)$ is the usual frequency-dependent environmental function given by [48, 5]
@@
When a voltage $V$ is applied to the free edge channel on the right, the tunneling current can be cast into the form  [^1] The charge current takes the standard form $I=|\gamma_{t}|^{2}\int^{\infty}_{-\infty}dt\;e^{iVt}\Big[g^{<}(t)G^{<}(t)-g^{>}(t)G^{>}(t)\Big]$.
@@
In the asymptotic limit of $T\ll E_{C}$, the equation (38) provides weakly temperature-dependent linear-response conductance $G(T)=\mathcal{F}G_{c}(T/E_{C})^{8}$, where $\mathcal{F}$ is a constant of order unity. The temperature scaling exponent is governed by four pairs of electron-hole excitations each contributing $T^{2}$ to the conductance, and thus the setup with large number of sites becomes essentially an electronic insulator at low temperature. All the previous treatments can be applied to compute the finite backscattering corrections to equation (38). Inclusion of such backscattering corrections will amount to the addition of linear in $|r_{u/v}|$ terms to the constant $\mathcal{F}$.
Figure 3: (a) Typical realization of circuits with two channels and two sites in a quantum Hall system. (b) Setup (a) with one of the channels looped. The red dot represents finite backscattering, while ithe cross symbol indicates that the corresponding QPC is fully opened. Note that the gates are not explicitly shown (see the text for details).
@@
VIII Multi-channel mesoscopic quantum Hall circuits
In the previous sections, we studied the QH circuits in which each grain accommodates a single channel to the left and also to the right of it. More generally, it is possible to engineer a mesoscopic QH circuit with multiple edge channels traversing each metal grain [23]. The fully ballistic regime of such multichannel QH circuits can be understood as follows. A portion of QH circuit outside the metal grain that consists of left and right moving edges can be described by a chiral bosonic field. Therefore, for the setup with $N$ identical grains, each hosting $M$ edge channels, there are in total $M(N+1)$ bosonic modes describing the corresponding plasmonic excitations. At energies much lower than the charging energy, and provided that the inter-grain Coulomb interactions are negligible, $N$ of the $M(N+1)$ bosonic modes are gapped by the charging effects and can be integrated out. The corresponding low-energy effective Hamiltonian is thus expressed in terms of the remaining $N_{\rm f}=\left[M(N+1)-N\right]$ gapless modes. Standard circuit analysis then shows that the conductance of each free channel takes the value $G_{0}=M/[2\pi(N+1)]$.
When $N_{\rm f}>1$, the problem can in principle be mapped onto that of a particle moving in a multidimensional periodic potential subject to dissipation [51, 52]. The usual perturbative calculations at large transparencies are generally unstable, in the sense that as temperature/voltage tends to zero, the perturbative corrections blow up. To explicitly demonstrate this behavior, we now study a typical multichannel QH circuit as schematically shown in Fig. 3(a). It consists of two islands and a total of six edge channels. For the clarity of presentation, we refer to the portion of the circuit to the left of the first island as the left part, that between the two grains as the middle part and that right of the second grain as the right part. Each part accommodates two channels (blue and black), here for generality, implemented via six QPCs as shown in Fig. 3(a).
@@ VIII Multi-channel mesoscopic quantum Hall circuits
IX Exotic quantum critical points in multi-channel QH circuits
As long as more than one gapless mode governs the low-energy dynamics, a mapping to the boundary sine-Gordon model is no longer applicable. Nonetheless, following Ref. [23], we show that in the multichannel QH circuits the number of gapless modes can be reduced to one by looping a selected number of edge channels. This method enables the realization of zero-temperature quantum critical phenomena characterized by a general scaling dimension $\eta<1/2$, where $\eta$ depends on both the number of grains and the number of looped edge channels.
To illustrate the looping method, we consider a simples setup as shown in Fig. 3(b), obtained by looping the second edge channel of Fig. 3(a) back onto the grains [54]. The key idea is that transport through the left and right QPCs (QPC 1 and 3 respectively) is unaffected by the presence of the looped part (on the left and right) of the circuit. Therefore, if both channels in the middle part are fully ballistic, they merely provide impedances to the incoming electrons at the left and right QPCs,and the entire setup becomes effectively equivalent to a single-grain device with two QPCs. The crucial difference here is that the number of connecting ballistic channels contributes an additional impedance, which in turn controls the scaling dimension of the boundary electron operators.
Following the equation-of-motion approach to account for the charging energy part of the Hamiltonian, one can express all bosonic fields in terms of incoming fields. The fields $\Phi_{\rm o}$ describing the two looped parts of the circuit, do not directly enter the transport problem through QPCs 1 and 3, and possess the correlator of the form
@@ IX Exotic quantum critical points in multi-channel QH circuits
with the gate voltage parameters $N_{\rm L,g}=2\pi\left(3N_{\rm 1g}+2N_{\rm 2g}\right)/5$ and $N_{\rm R,g}=2\pi\left(2N_{\rm 1g}+3N_{\rm 2g}\right)/5$. The boundary term (74) having scaling dimensions $2/5$ supports a zero-temperature quantum critical point that can be reached by varying gate voltages and/or backscattering amplitudes. In addition, here the boundary sine-Gordon model can be exploited to uncover all transport and thermodynamic properties. In the case of $N$ grains and $M$ looped channels, see Fig. 3(b) for $N=2$ and $M=1$, the backscattering in the left and right QPCs can be described by a single gapless field, and the leading-order perturbations have the scaling dimension $\nu_{\rm o}=(M+1)/(2M+N+1)$. All these different realizations support distinct zero-temperature quantum critical points, where for $\nu_{\rm o}<1/4$, the higher order backscattering terms also have to be taken into account. Therefore, channel looping in multi-site QH circuits provides a powerful and experimentally accessible route to engineer exotic quantum critical points, yielding far-reaching insights into non-Fermi liquid behavior.
X Heating effects in a multi-site QH circuit
In section VI, we presented results for the transport properties of the four-site QH circuit in a fully non-equilibrium regime driven by an voltage bias. In that analysis, we neglected the effects coming from the heating of the grains by injected voltage, namely the Joule heat $J_{V}=V^{2}/2R_{q}$ carried by each incoming chiral edge due to the corresponding voltage drop $V$. Likewise, at a given temperature $T$, the heat current carried by a free chiral edge described by the bosonic field $\Phi^{0}$ takes the form